please help me solve logx square root 3/3= -1/2
Answer by lwsshak3(11628)

solve logx square root 3/3= -1/2
log(x√3/3)=-1/2
10^(-1/2)=x√3/3
0.3162=x√3/3
√3x=3*.3162=0.9487
x=0.9487/√3=0.5477